OpenStudy (anonymous):

Verify: cos (x- π/4) = √2/2(cosx+sinx) Need help ASAP pleasee

OpenStudy (anonymous):

can i get the answer choices because im not sure if its true or false or number or equation because if its true or false then its true

OpenStudy (anonymous):

It isn't multiple choose question I have to show proof like how they equal to eachother

OpenStudy (tkhunny):

Why not just verify it? \(\cos(\alpha - \beta) = \cos(\alpha)\cos(\beta) + \sin(\alpha)\sin(\beta)\)

OpenStudy (anonymous):

I am supposed to verify it but what your saying is that if I use cos (a-b) I will be able to prove it?

OpenStudy (tkhunny):

Yup. You'll be done.

OpenStudy (anonymous):

Cool thank you!